OpenStudy (melissa_something):

MEDAL <3 Log question: Evaluate with out calculator:

OpenStudy (melissa_something):

\[\log _{64}4= ?\]

OpenStudy (melissa_something):

Ik the answer is 1/3 but I don't know how or why

OpenStudy (amistre64):

lets say y = log_64 (4) what would you say 'undoes' the log? any thoughts?

OpenStudy (melissa_something):

So b raised to ___ =64?

OpenStudy (melissa_something):

Ik they're similar in some way I just don't see the connection :/

OpenStudy (amistre64):

recall the log rule: sounds like your thinking of the righ tthing b^(log_b (x)) = x so lets base 64 both sides to undo th elog_64 64^y = 4 now i spose the way i would attempt a solution is to see if 4 and 64 have any factors in common

OpenStudy (melissa_something):

uhhh ok let me write this :)

OpenStudy (melissa_something):

Ah yay, factoring -.-

OpenStudy (amistre64):

what is 64/4?

OpenStudy (melissa_something):

16!

OpenStudy (melissa_something):

Omg wait 3 times 4 is 16

OpenStudy (melissa_something):

Is it like that for all of them?

OpenStudy (melissa_something):

No it's not I just checked on another problem -___- oh well

OpenStudy (amistre64):

so: 4*16 = 64, or 4*4*4 = 64 lets cube both sides .... \[(64^y)^3=(4)^3\] \[(64^y)^3=(64)^1\] now comes another rule: \[\large (a^m)^n=a^{mn}\] apply it \[(64^{3y}=(64)^1\] what can we determine now?

OpenStudy (melissa_something):

My brain would have never thought like that... Is there any tips you have because I would've never known 4*4*4=64?

OpenStudy (amistre64):

when you hit a road block, think of ways to get past it. i know that if this has any possible way of being 1/3 that both sides must have something in common. 64 and 4 would have to be related in some fashion. common factors seemed appropriate to me.

OpenStudy (amistre64):

64^(3y) = 64^(1) if and only if ... 3y=1 what is y?

OpenStudy (melissa_something):

Ok ok let me process this! & give u a medal before I forget but plz wait D:

OpenStudy (melissa_something):

Uh, I'm not sure why 3y has to = 1?

OpenStudy (melissa_something):

OH I GET IT THE BASES ARE THE SAME!

OpenStudy (amistre64):

summary \[y=log_{64}(4)\] \[64^y=64^{log_{64}(4)}\] \[64^y=4\] \[note:4^3=64~:~so~cube~it~all\] \[(64^y)^3=(4)^3\] \[rule:(a^m)^n=a^{mn}~:~use~it\] \[64^{(3y)}=64^{(1)}\] \[3y=1\]

OpenStudy (amistre64):

yes, same bases :)

OpenStudy (amistre64):

if we log_64 both sides the 64s clear out and we are left with 3t=1 if you want to include a step

OpenStudy (melissa_something):

Wow thank you, you didn't have to!!

OpenStudy (amistre64):

youre welcome, and good luck ;)

OpenStudy (melissa_something):

Okay, so we just- OMG I get it thank you!!!!

OpenStudy (melissa_something):

Two questions, how long did it take you to get to 99?

OpenStudy (amistre64):

i was at the top long before they started using number values .. so i was 99 when they started this new system by default.

OpenStudy (melissa_something):

Also um.... uhhh.. I forgot damn it. Oh well, thank you!

OpenStudy (melissa_something):

OH WAIT I REMEMBER: Is there a way to view all your past asked questions?

OpenStudy (amistre64):

placement was based on fan counts when i started. then they applied silly titles, then a medal system was integrated, ive been pretty much just floating by at the top ....

OpenStudy (melissa_something):

Ok thanks got it haha! :D It would be cool if they gave scholarships for doing this!

OpenStudy (amistre64):

hover your icon in the top right corner and go to your 'profile' page. questons asked and answered is an option after it loads.